OpenStudy (anonymous):

Solve the problem as directed. If x varies inversely as y, and x = 11 when y = 15, find x when y = 3. x=?

OpenStudy (anonymous):

55

OpenStudy (anonymous):

x=k/y 11=k/15 k=165 x=165/3 x=55

OpenStudy (anonymous):

thank you !